OpenStudy (anonymous):

Divide & express result in standard form... 8 over 5-i

OpenStudy (anonymous):

\[\frac{8}{5-i}=\frac{8}{5-i}\times \frac{5+i}{5+i}=\frac{8(5+i)}{5^2+1^2}\] \[=\frac{40+8i}{26}\] then standard form

OpenStudy (anonymous):

its asking for two fractions...